OpenStudy (anonymous):

y=75x+175? trying to find x-intercept

OpenStudy (anonymous):

75x+175=0 75x=-175 x=-2 1/3

jimthompson5910 (jim_thompson5910):

To find the x intercept, plug in y = 0 and solve for x y=75x+175 0=75x+175 -175 = 75x -175/75 = x -7/3 = x x = -7/3 So the x intercept is (-7/3, 0)

OpenStudy (anonymous):

so in decimal form? for -7/3? so i can plot.

jimthompson5910 (jim_thompson5910):

-7/3 = -2.333333... (the 3s repeat forever)

jimthompson5910 (jim_thompson5910):

so x intercept is (-2.3333, 0)

OpenStudy (anonymous):

what about -2.3, 0

jimthompson5910 (jim_thompson5910):

sure that'll work too